Assume Kader Company has the following reported amounts: Sales revenue $510,000

Brief Exercise 5-10 

Assume Kader Company has the following reported amounts: Sales revenue $510,000, Sales returns and allowances $15,000, Cost of goods sold $330,000, and Operating expenses $110,000. 

(a) Compute net sales. 

Net sales = $_______ 

(b) Compute gross profit. 

Gross profit  = $_______ 

(c) Compute income from operations. 

Income from operations  = $_______

(d) Compute the gross profit rate. (Round answer to 1 decimal place, e.g. 25.29.) 

Gross profit rate = $_______
Answer #1

Net Sales = Sales - returns and allowances

= 510,000-15,000

= 495,000

Gross profit = Net sales - Cost of Goods Sold

= 495,000-330,000

= 165,000

Income from operations = Gross Profit - Operating expenses

= 165,000-110,000

= 55,000

Gross profit rate = Gross Profit/net sales

= 165,000/495,000

= 33.3%
Answer #2
Net income
answered by: Delwar Hussain